OpenStudy (anonymous):

y=(1-x)/(1+x) at (-2,-3) i suppose to find equation at tangent line the answer i got is y=-2x-7

OpenStudy (idku):

Sure, did you get the derivative? Don't use quotient tho

OpenStudy (anonymous):

but my professor answers are all positive

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

his answer is 2x+y+7=0

OpenStudy (idku):

(1-x)/(1+x) = (-x+1)/(x+1)= (-x-1+2)/(x+1)=(-x-1)/(x+1) + 2/(x+1) = -1 + 2/(x+1)

OpenStudy (anonymous):

i'm thinking he wrote it in implicit form and changed everything arounf

hartnn (hartnn):

y = -2x-7 is same as 2x+y+7 = 0

hartnn (hartnn):

so you and your prof, both are correct

OpenStudy (anonymous):

ok